Shevlin Inc. had 700,000 shares of common stock outstanding at December 31, 2019. Shevlin issued an additional 60,000 shares of common stock on October 1, 2020. Shevlin also had $20,000,000 of 3% convertible bonds outstanding throughout 2020, which are convertible into a total of 400,000 shares of common stock. These convertible bonds were originally issued at par. Shevlin’s net income for the year ended December 31, 2020, was $2,200,000. Assuming the applicable income tax rate was 25%, diluted earnings per share for 2020 should be

A. $1.97 per share

B. $2.00 per share

C. $2.11 per share

D. $2.37 per share

E. None of above

Please show your steps!
Answer #1

Solution:

Weighted average outstanding shares in 2020 = 700000 + (60000*3/12) = 715000 shares

Net Income for 2020 = $2,200,000

Interest expense on bond (Net of tax) = $20,000,000*3% (1-0.25) = $450,000

Future net income after conversion of bond = $2,200,000 + $450,000 = $2,650,000

Future weighted average outstanding shares = 715000 + 400000 = 1115000 shares

Diluted Earning per share for 2020 = Future net income / Future weighted average outstanding shares

= $2,650,000 / 1115000 = $2.37 per share

Hence option D is correct.